OpenStudy (anonymous):

Use basic identities to simplify the expression. csc(theta) * cot(theta) / sec (theta)

OpenStudy (anonymous):

I got really stuck on this one

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ \sin (\theta) }\frac{ \cos (\theta) }{ \sin(\theta) } }{ \frac{ 1 }{ \cos(\theta)} }\] \[\frac{ \frac{ \cos(\theta) }{ \sin ^{2}(\theta) } }{ \frac{ 1 }{ \cos(\theta) } }\] \[\frac{ \cos(\theta) }{ \sin ^{2}(\theta) }(\frac{ \cos \theta) }{ 1 })\] \[\frac{ \cos ^{2(\theta)} }{ \sin ^{2}(\theta) }\] \[\tan ^{2}(\theta)\]

OpenStudy (anonymous):

not tan^2(theta) \[\cot ^{2}(\theta)\]

OpenStudy (anonymous):

thank you but that is not one of my options from my homework. It can be cot^2, csc^2, sec^2 or 1.

OpenStudy (anonymous):

oh ok! Thank you!!!